# Microscopic Mechanism of Anyon Superconductivity Emerging from Fractional Chern Insulators

### v1

@@ III.2 Equivalent picture using fermionic partons
Our parton construction in the previous Eq. (15) was chosen to be able to conveniently describe the FQH and semion crystal phases at different fillings, as well as the superconductors that emerge on doping them. Here, we outline how some of the same results at particular fillings can be recovered using just the fermionic partons and dispensing with the boson. This specific parton ansatz was used in [23] to analyze the doped 2/3 fractional Chern insulator; here we adapt it to our setting and supply an independent derivation highlighting how a charge-2 superconductor emerges upon doping, generalizing the approach to generic spin singlet FQH states at filling $\nu=2/(2n\pm 1)$; see Appendix A.
Before delving into finite doping, let us briefly specialize to $\nu=2/3$ and discuss how all three phases, FQH, semion crystal and superconductor can be accessed at fixed density within the fermionic parton theory $e_{\sigma}(x)=f_{\sigma}g_{1}g_{2}$ While energetics may make a superconductor at the same filling unlikely, an effective theory will help us build intuition for the parton description. Indeed, starting from the parton Chern number assignments $(-2,1,1)$ for the $(f_{\sigma},\,g_{1},\,g_{2})$ partons, it is easily seen that the semion crystal is accessed with the assignments $(-2,1,0)$ while the superconductor with $(-2,1,-2)$. Since this only involves changing the topology of the last parton, that can be modeled as sign changes of Dirac fermion mass, which models the topology-changing transitions. Moreover, the magnetic translation symmetry resulting from the fact that $g_{2}$ sees emergent gauge flux of $1/3$, implies three valleys related by translations. This gives rise to three species of low-energy fermions $\Psi_{i}$, which are related to one another by magnetic translations. Ultimately, we can write down the following effective theory:
@@
Density tuned transitions from fermionic partons.— For simplicity, we consider particle doping $\delta>0$. To form a superconductor upon doping, the Chern number of one of the auxiliary partons, say $g_{2}$, must change from $1$ to $-2$. To keep both the spinons $f_{\sigma}$, and the other chargon $g_{1}$ at a fixed integer levels $(-2,1)$ requires the emergent gauge fields to adjust such that $g_{2}$ experiences a magnetic flux of $n_{\Phi}^{g_{2}}=\Phi_{g_{2}}/\Phi_{0}=-\frac{1}{3}-\frac{\delta}{2}$, while being at a density of $n_{g_{2}}=\frac{2}{3}+\delta$. In the presence of a lattice, the total Chern number of $g_{2}$ is constrained by a Diophantine equation [44, 45, 46]:
@@
For Eq. (39) to be satisfied, terms involving the doping $\delta$ must cancel while also recovering $C_{g_{2}}=1$ in the limit $\delta\rightarrow 0$. Thus, we find $C_{g_{2}}=-2$ as a consistent solution in the doped case (see Appendix D for more details), which signals the formation of charge-2 superconductivity. This illustrates, why the underlying lattice potential is crucial for establishing superconductivity upon doping the FCI. For the FCI at $2/3$ filling, this establishes a relation to previous work [23], which we obtain in an independent approach. Performing a continuum expansion and taking into account that due to the flux seen by $g_{1,2}$, there are three equivalent minima $\mathbf{K}_{i}$, we find
@@
Upon doping, the finite spin gap close to the critical point ensures that the spinons remain in their gapped IQH state. Consequently, it is energetically favorable to dope spin-singlet anyons which carry charge $2/3$ and correspond to $g_{1,2}$. When changing the electron density by $\delta$, we need to adjust the densities of all partons $f_{\sigma},g_{1},g_{2}$. To ensure that the spinons remain in a gapped level $-2$ IQH state, we must therefore change the flux they experience. Specifically, their flux changes by $-\delta/2$ according to the Streda formula. Since the partons $g_{1,2}$ are coupled to the spinons through the gauge fields $a_{1,2}$,
@@
where $\mathbf{a}_{\delta}$ is a gauge field generating flux $\Delta\Phi_{g_{2}}$. Consequently, Landau levels form in the doped valleys, where the cyclotron frequency scales with doping $\Delta\Phi_{g_{2}}\propto-\delta$. This suggests that the stiffness of the resulting superconductor will scale with doping. With this flux assignment for $a_{1/2}$, parton $g_{1}$ remains in a $C=1$ band after doping, while $g_{2}$ fills an integer number of Landau levels in the doped valleys, such that its total Chern number is $C=-2$; see Table 2 for an overview. This can be seen as follows: After doping, the density for $g_{2}$ is $n_{g_{2}}=2/3+\delta$ per unit cell, and it feels a flux of $\Phi_{g_{2}}=1/3-\delta/2$ in units of $\Phi_{0}$. On the FQH side of the transition, the filled band has $C=1$. The Streda formula implies that a change in the flux leads to a change in the density of partons that fill that band. Hence, we find that only $n_{C=1}=2/3-\delta/2$ partons fit in the originally filled band, leaving a density of $\Delta n=n_{g}-n_{C=1}=3\delta/2$ to fill the valleys in the second band. When the valley symmetry of the partons is preserved, the doped partons fill $\nu=\Delta n/(3\Delta\Phi)=-1$ Landau levels in all three valleys. Alternatively, if translations are broken, either explicitly or spontaneously upon doping, valley symmetry is broken. In this case, all doped partons enter into a single valley, where they fill $\nu=\Delta n/\Delta\Phi=-3$ Landau levels; see Fig. 3. Doping in a valley-broken state corresponds to the vortices forming a Jain state, Eq. (29), while doping the partons in a valley-symmetric state corresponds to the vortices forming a valley singlet, Eq. (30).
@@
The final Chern number for the $g_{2}$ partons is the sum of the originally filled lower Chern band and the Landau levels in the doped valley, giving $C=-2$ in total, on both sides of the transition, as suggested by the Diophantine equation Eq. (39). This highlights the importance of a lattice or at least a periodic potential, as the same arguments cannot be applied to a system with continuous translations in flat Landau levels.
@@
which leaves us with the same theory, Eq. (35), as in Sec. III.1, describing a charge-2e superconductor, without residual topological order and a spin quantum Hall response of $-2$. Finally, we argue that the doped state has zero electrical resistivity, as required for a superconductor. To that end, note that gluing the partons $g_{1}$ and $g_{2}$ at level (1,-2) back together after doping, can be viewed as forming a bosonic IQH state for the composite boson $g_{1}g_{2}$, characterized by a Hall conductivity of $\sigma_{xy}=+2$ [47, 48, 43]. This exactly cancels the Hall resistivity of the spinons, resulting in a vanishing total resistivity according to the Ioffe-Larkin rule: $\rho_{e}=\rho_{b}+\rho_{f}=0$ [42].
@@
$f_{\sigma}$ $\frac{1}{6}+\frac{\delta}{2}$ $-(\frac{1}{6}+\frac{\delta}{2})\Phi_{0}$ $-2\rightarrow-2$
$g_{1}$ $\frac{1}{3}+\delta$ $(\frac{1}{3}+\delta)\Phi_{0}$ $+1\rightarrow+1$
$g_{2}$ $\frac{1}{3}+\delta$ $(\frac{1}{3}-\frac{\delta}{2})\Phi_{0}$ $+1\rightarrow-2$
@@
Doping the $\nu=2/5$ state.— In the $(332)$ state, the spinons have total Chern number $C=+2$. Applying similar reasoning as for the $(\bar{1}\bar{1}2)$ state, adjusting the flux for the spinons such that they remain in their gapped $C=2$ state, introduces a change in flux for the partons $g_{1/2}$, as outlined in Table 3. Assuming that all doped partons enter into a unique valley, which implies translation symmetry breaking we find that $g_{1}$ has total Chern number $C=-1$ after doping, and $g_{2}$ has $C=2$. The resulting superconductor is the same as for $\nu=2/3$, Eq. (35), but with a spin Hall response of $+2$ instead of $-2$. While the superconductor, with broken valley symmetry for the partons, corresponding to broken translations for the electronic state, straightforwardly generalizes from $\nu=2/3$ to $\nu=2/5$, the valley-symmetric state of $\nu=2/3$ does not have a counterpart at $\nu=2/5$, where such valley symmetric states map generically to more exotic superconductors. From this perspective, the $\nu=2/3$ state is special, as here the valley-symmetric and valley-broken superconductors are topologically equivalent.
@@
$f_{\sigma}$ $\frac{1}{10}+\frac{\delta}{2}$ $(\frac{1}{10}+\frac{\delta}{2})\Phi_{0}$ $+2\rightarrow+2$
$g_{1}$ $\frac{1}{5}+\delta$ $(\frac{1}{5}-\delta)\Phi_{0}$ $+1\rightarrow-1$
$g_{2}$ $\frac{1}{5}+\delta$ $(\frac{1}{5}+\frac{\delta}{2})\Phi_{0}$ $+1\rightarrow+2$
@@ III.3 Doping the semion crystal
where we have explicitly reinstated a charge $-e$ bosonic field $\varphi$, corresponding to the chargon in the parton decomposition $e_{\sigma}=f_{\sigma}b^{\prime}$, which is in a translation symmetry-breaking Mott phase but close to its critical point. Doping this chiral spin-liquid works similarly to previous proposals [42, 27], with the difference of translational symmetry breaking in the semion crystal we discuss. In a regular Kalmeyer-Laughlin spin liquid, the charge gap is much larger than the spin gap $\Delta_{c}\gg\Delta_{s}$, which implies that the lowest lying anyons carry spin 1/2. In stark contrast, the lowest energy excitations of the semion crystal close to the QCP are spinless, charge $Q=-e$ semions described by $\varphi(x)$. Since $\Delta_{c}/\Delta_{s}\rightarrow 0$ it is natural to think of these anyons as fractionalized Cooper pairs. As Sec. II.3 shows, a similar interpretation is possible on the FCI side of the transition.
Upon hole doping, the density of doped chargons is $-\delta$. To keep the spinons gapped, we must adjust the flux $\frac{1}{2\pi}\nabla\times\alpha\sim\mp\delta/2$, leading to a filling $\nu_{b}=\pm 2$. Consequently, the $b$ bosons form a bosonic integer quantum Hall state, which when integrated out precisely cancels the topological order in (49) arising from the spinons. We expect that the translational symmetry breaking is preserved upon doping, which leads to a charge-2 SC without residual topological order in the presence of a charge density wave, see Fig. 1(b).
@@ III.4 Properties of the anyon superconductor
At $2/3$ filling, both options we discussed are topologically ‘d-id’ superconductors. Owing to the fact that the neutral spin modes on the boundary are counter-propagating and remain unmodified close to the critical point, the superconducting state retains a spin response $-\frac{2}{4\pi}A_{s}\wedge\mathrm{d}A_{s}$ which is opposite to the electronic Hall effect. Due to the nature of the FQH parent states at general fillings $\nu=2/(2l+1)$, the sign of the spin-quantum Hall effect (and therefore also the thermal Hall response) alternates as $\sigma_{xy}^{s}=2\cdot(-1)^{l}$
@@
As we are describing an insulator-to-insulator transition, this behavior is expected. Remarkably, since our doping procedure preserves the spin-gap, we expect that it also sets the size of Cooper pairs for weak doping. This implies that the superconductors we describe are inherently in a Bose-Einstein Condensate limit and the critical temperature is set by the stiffness, which grows with the doping level.
@@
We note that charge-2 superconductivity with opposing thermal Hall response can be obtained by considering Kohn-Luttinger mechanisms in the complementary weakly interacting limit where the parent state is a chiral metal [49, 50].
@@ IV Microscopic lattice model
where $t$ ($t^{\prime}$) is the (next-)nearest neighbor hopping, $U$ are interlayer interactions and $V$ nearest-neighbor interactions both within and between the layers. The Peierls phases are chosen to match the flux through a triangle $\Phi=B_{z}a^{2}\sqrt{3}/4$, where $a$ is the lattice constant. At $\Phi=\frac{\Phi_{0}}{2}$ flux piercing a triangle, and $t^{\prime}=t/4$ the system exhibits nearly flat particle-hole symmetric valence and conduction bands, with Chern numbers $C=1$ and $C=-1$, respectively; see Fig. 4(a). Our model explicitly breaks time-reversal symmetry, and the Chern numbers for the two flavors $\sigma$ are equal. At finite filling and intermediate interaction strength, the system can thus realize various bilayer quantum Hall states. The Hubbard-Hofstadter model discussed here should be thought of as a controlled system, with potential experimental implementation in moiré heterostructures [29]. It serves as proof of principle that the considerations of the previous sections can be realized in a microscopic lattice model, but we expect our mechanism to be applicable more generally.
@@
First, we establish that the Halperin $(\bar{1}\bar{1}2)$ state is stabilized in the bilayer Hubbard-Hofstadter model Eq. (51) on the triangular lattice. In contrast to the $(110)$ state discussed in Refs. [29, 27], the $(\bar{1}\bar{1}2)$ state has a non-trivial topological order and has fascinating properties in its own right. As discussed in Sec. II.1, it can be understood in a parton language as spinons filling Chern bands with $C=-1$ and chargons in a bosonic $1/2$ Laughlin state. We uniquely identify the $(\bar{1}\bar{1}2)$ state from its entanglement spectrum and by performing adiabatic flux insertion into the cylinder.
@@
±$\pm\partial_{t}\varphi_{s}\partial_{y}\varphi_{s}+v_{s}\partial_{y}\varphi_{s}\partial_{y}\varphi_{s}]$
for $(mmn)$ states with $m=n\pm 1$ and velocities $v_{c/s}>0$. Note that for $m=n+1$, the two edge modes have the same chirality, while for $m=n-1$, they have opposite chirality, resulting in a non-chiral entanglement spectrum, assuming $n>0$. We find the edge Hamiltonian and total momentum [53]
@@
Using the above, the edge spectrum accessible through the entanglement spectrum can be understood, see Appendix E for details. There are three dispersing modes: A pure charge mode with $L_{s}=0$, a pure spin mode with $L_{c}=0$, and a mixed mode with $L_{c}\neq 0\neq L_{s}$. The counting of the entanglement spectrum in Fig. 5 agrees well with the expected edge spectrum for the $(\bar{1}\bar{1}2)$ state for which $m=1,\,n=2$, as graphically illustrated.
@@
Increasing the nearest-neighbor interactions triggers a phase transition and tunes the system into a different insulator, which we will argue is nothing but the semion crystal. While the charge response of this phase is trivial, spin-pumping reveals a quantized spin-Hall response, which is identical to the spin-Hall response of the $(\bar{1}\bar{1}2)$ state; see Fig. 6(b). For general semion crystals formed at filling $\nu=2/(2n\pm 1)$, the quantum spin Hall response is alternating $Q_{s}=\pm 1$. This can be understood in this limit as lattice translations are broken, and the unit cell is enlarged. Within the enlarged unit cell, the electron and spinon density is integer $(n_{e}=1)$, while the flux per unit cell of the crystal is ±$\pm\Phi_{0}/2$ modulo $2\pi$. Thus, the spinon density in the crystal is now large enough to absorb all of the external flux by forming an IQH state at level ±$\pm 2$, while the chargons are in a Mott state with respect to the enlarged unit cell. As a concrete example, consider the 2/3 FCI, which transitions into the semion crystal, shown in the top left of the phase diagram of Figure 4(b). While triangles of the lattice enclose $\pi/2$ flux, the larger triangles defined by the charge ordered state have three times the area and enclose a flux of $23\pi/2\equiv-\pi/$ 2, which reverses the sign of the spin Hall conductance compared to the chiral spin liquid at integer filling.
Figure 7: Evidence for second-order phase transition in the Hubbard-Hofstadter model. (a) Order parameter $\Delta_{\text{CDW}}$ as a function of the interaction strength $V$ for fixed $U=4t$, $\chi=2048$ and $L_{y}=4$. Inset: Density correlations $C(k_{y}=0,x)=\langle\rho_{-k_{y}}(x)\rho_{k_{y}}(0)\rangle$ for $V=1.1t$ and $V=2.0t$. Due to the finite cylinder circumference, translations are already weakly broken in the FQH phase. (b) The correlation length of the density mode, i.e. of operators with zero charge and spin $(Q,S_{z})=(0,0)$, peaks at a certain interaction strength $V$ which we identify as the critical point (indicated by a dashed line in all panels). By contrast, the spin correlation length $(Q,S_{z})=(0,\hbar)$ remains small and does not grow when approaching the transition. (c) The ratio of $\xi_{2e}$ [$(Q,S_{z})=(2,0)$] to $\xi_{1e}$ [$(Q,S_{z})=(1,\hbar/2)$] correlation length is much larger than the trivially expected value of $0.5$ and peaks close to the critical point, growing as a function of the bond dimension $\chi$.

### v2

@@ III.2 Equivalent picture using fermionic partons
Our parton construction in the previous Eq. (15) was chosen to be able to conveniently describe the FQH and semion crystal phases at different fillings, as well as the superconductors that emerge on doping them. Here, we outline how some of the same results at particular fillings can be recovered using just the fermionic partons and dispensing with the boson. This specific parton ansatz was used in [67] to analyze the doped 2/3 fractional Chern insulator; here we adapt it to our setting and supply an independent derivation highlighting how a charge-2e superconductor emerges upon doping, generalizing the approach to generic spin singlet FQH states at filling $\nu=2/(2n\pm 1)$; see Appendix A.
Before delving into finite doping, let us briefly specialize to $\nu=2/3$ and discuss how all three phases, FQH, semion crystal and superconductor can be accessed at fixed density within the fermionic parton theory $e_{\sigma}(x)=f_{\sigma}g_{1}g_{2}$ . While energetics may make a superconductor at the same filling unlikely, an effective theory will help us build intuition for the parton description. Indeed, starting from the parton Chern number assignments (2,1,1)(-2,1,1) for the $(f_{\sigma},\,g_{1},\,g_{2})$ partons, it is easily seen that the semion crystal is accessed with the assignments (2,1,0)(-2,1,0) while the superconductor with (2,1,2)(-2,1,-2). Since this only involves changing the topology of the last parton, that can be modeled as sign changes of Dirac fermion mass, which models the topology-changing transitions. Moreover, the magnetic translation symmetry resulting from the fact that $g_{2}$ sees emergent gauge flux of $1/3$, implies three valleys related by translations. This gives rise to three species of low-energy fermions $\Psi_{i}$, which are related to one another by magnetic translations. Ultimately, we can write down the following effective theory:
@@
Density tuned transitions from fermionic partons.— For simplicity, we consider particle doping $\delta>0$. To form a superconductor upon doping, the Chern number of one of the auxiliary partons, say $g_{2}$, must change from $1$ to 2-2. To keep both the spinons $f_{\sigma}$, and the other chargon $g_{1}$ at a fixed integer levels (2,1)(-2,1) requires the emergent gauge fields to adjust such that $g_{2}$ experiences a magnetic flux of $n_{\Phi}^{g_{2}}=\Phi_{g_{2}}/\Phi_{0}=-\frac{1}{3}-\frac{\delta}{2}$, while being at a density of $n_{g_{2}}=\frac{2}{3}+\delta$. In the presence of a lattice, the total Chern number of $g_{2}$ is constrained by a Diophantine equation [76, 7, 1]:
@@
For Eq. (39) to be satisfied, terms involving the doping $\delta$ must cancel while also recovering $C_{g_{2}}=1$ in the limit $\delta\rightarrow 0$. Thus, we find $C_{g_{2}}=-2$ as a consistent solution in the doped case (see Appendix D for more details), which signals the formation of charge-2e superconductivity. This illustrates, why the underlying lattice potential is crucial for establishing superconductivity upon doping the FCI. For the FCI at $2/3$ filling, this establishes a relation to previous work [67], which we obtain in an independent approach. Performing a continuum expansion and taking into account that due to the flux seen by $g_{1,2}$, there are three equivalent minima $\mathbf{K}_{i}$, we find
@@
Upon doping, the finite spin gap close to the critical point ensures that the spinons remain in their gapped IQH state. Consequently, it is energetically favorable to dope spin-singlet anyons which carry charge $2/3$ and correspond to $g_{1,2}$. When changing the electron density by $\delta$, we need to adjust the densities of all partons $f_{\sigma},g_{1},g_{2}$. To ensure that the spinons remain in a gapped level 2-2 IQH state, we must therefore change the flux they experience. Specifically, their flux changes by $\delta/2$ according to the Streda formula. Since the partons $g_{1,2}$ are coupled to the spinons through the gauge fields $a_{1,2}$,
@@
where $\mathbf{a}_{\delta}$ is a gauge field generating flux $\Delta\Phi_{g_{2}}$. Consequently, Landau levels form in the doped valleys, where the cyclotron frequency scales with doping $\Delta\Phi_{g_{2}}\propto-\delta$. This suggests that the stiffness of the resulting superconductor will scale with doping. With this flux assignment for $a_{1/2}$, parton $g_{1}$ remains in a $C=1$ band after doping, while $g_{2}$ fills an integer number of Landau levels in the doped valleys, such that its total Chern number is C=2C=-2; see Table 2 for an overview. This can be seen as follows: After doping, the density for $g_{2}$ is $n_{g_{2}}=2/3+\delta$ per unit cell, and it feels a flux of $\Phi_{g_{2}}=1/3-\delta/2$ in units of $\Phi_{0}$. On the FQH side of the transition, the filled band has $C=1$. The Streda formula implies that a change in the flux leads to a change in the density of partons that fill that band. Hence, we find that only $n_{C=1}=2/3-\delta/2$ partons fit in the originally filled band, leaving a density of $\Delta n=n_{g}-n_{C=1}=3\delta/2$ to fill the valleys in the second band. When the valley symmetry of the partons is preserved, the doped partons fill $\nu=\Delta n/(3\Delta\Phi)=-1$ Landau levels in all three valleys. Alternatively, if translations are broken, either explicitly or spontaneously upon doping, valley symmetry is broken. In this case, all doped partons enter into a single valley, where they fill $\nu=\Delta n/\Delta\Phi=-3$ Landau levels; see Fig. 3. Doping in a valley-broken state corresponds to the vortices forming a Jain state, Eq. (29), while doping the partons in a valley-symmetric state corresponds to the vortices forming a valley singlet, Eq. (30).
@@
The final Chern number for the $g_{2}$ partons is the sum of the originally filled lower Chern band and the Landau levels in the doped valley, giving C=2C=-2 in total, on both sides of the transition, as suggested by the Diophantine equation Eq. (39). This highlights the importance of a lattice or at least a periodic potential, as the same arguments cannot be applied to a system with continuous translations in flat Landau levels.
@@
which leaves us with the same theory, Eq. (35), as in Sec. III.1, describing a charge-2e superconductor, without residual topological order and a spin quantum Hall response of 2-2. Finally, we argue that the doped state has zero electrical resistivity, as required for a superconductor. To that end, note that gluing the partons $g_{1}$ and $g_{2}$ at level (1,-2) back together after doping, can be viewed as forming a bosonic IQH state for the composite boson $g_{1}g_{2}$, characterized by a Hall conductivity of $\sigma_{xy}=+2$ [45, 63, 18]. This exactly cancels the Hall resistivity of the spinons, resulting in a vanishing total resistivity according to the Ioffe-Larkin rule: $\rho_{e}=\rho_{b}+\rho_{f}=0$ [70].
@@
$f_{\sigma}$ $\frac{1}{6}+\frac{\delta}{2}$ $(\frac{1}{6}+\frac{\delta}{2})\Phi_{0}$ $2\rightarrow-2$
$g_{1}$ $\frac{1}{3}+\delta$ $(\frac{1}{3}+\delta)\Phi_{0}$ $1\rightarrow+1$
$g_{2}$ $\frac{1}{3}+\delta$ $(\frac{1}{3}-\frac{\delta}{2})\Phi_{0}$ $1\rightarrow-2$
@@
Doping the $\nu=2/5$ state.— In the $(332)$ state, the spinons have total Chern number $C=+2$. Applying similar reasoning as for the $(\bar{1}\bar{1}2)$ state, adjusting the flux for the spinons such that they remain in their gapped $C=2$ state, introduces a change in flux for the partons $g_{1/2}$, as outlined in Table 3. Assuming that all doped partons enter into a unique valley, which implies translation symmetry breaking we find that $g_{1}$ has total Chern number C=1C=-1 after doping, and $g_{2}$ has $C=2$. The resulting superconductor is the same as for $\nu=2/3$, Eq. (35), but with a spin Hall response of +2+2 instead of 2-2. While the superconductor, with broken valley symmetry for the partons, corresponding to broken translations for the electronic state, straightforwardly generalizes from $\nu=2/3$ to $\nu=2/5$, the valley-symmetric state of $\nu=2/3$ does not have a counterpart at $\nu=2/5$, where such valley symmetric states map generically to more exotic superconductors. From this perspective, the $\nu=2/3$ state is special, as here the valley-symmetric and valley-broken superconductors are topologically equivalent.
@@
$f_{\sigma}$ $\frac{1}{10}+\frac{\delta}{2}$ $(\frac{1}{10}+\frac{\delta}{2})\Phi_{0}$ $2\rightarrow+2$
$g_{1}$ $\frac{1}{5}+\delta$ $(\frac{1}{5}-\delta)\Phi_{0}$ $1\rightarrow-1$
$g_{2}$ $\frac{1}{5}+\delta$ $(\frac{1}{5}+\frac{\delta}{2})\Phi_{0}$ $1\rightarrow+2$
@@ III.3 Doping the semion crystal
where we have explicitly reinstated a charge e-e bosonic field $\varphi$, corresponding to the chargon in the parton decomposition $e_{\sigma}=f_{\sigma}b^{\prime}$, which is in a translation symmetry-breaking Mott phase but close to its critical point. Doping this chiral spin-liquid works similarly to previous proposals [70, 9], with the difference of translational symmetry breaking in the semion crystal we discuss. In a regular Kalmeyer-Laughlin spin liquid, the charge gap is much larger than the spin gap $\Delta_{c}\gg\Delta_{s}$, which implies that the lowest lying anyons carry spin 1/2. In stark contrast, the lowest energy excitations of the semion crystal close to the QCP are spinless, charge Q=eQ=-e semions described by $\varphi(x)$. Since $\Delta_{c}/\Delta_{s}\rightarrow 0$ it is natural to think of these anyons as fractionalized Cooper pairs. As Sec. II.3 shows, a similar interpretation is possible on the FCI side of the transition.
Upon hole doping, the density of doped chargons is $\delta$. To keep the spinons gapped, we must adjust the flux $\frac{1}{2\pi}\nabla\times\alpha\sim\mp\delta/2$, leading to a filling $\nu_{b}=\pm 2$. Consequently, the $b$ bosons form a bosonic integer quantum Hall state, which when integrated out precisely cancels the topological order in (49) arising from the spinons. We expect that the translational symmetry breaking is preserved upon doping, which leads to a charge-2e superconductor without residual topological order in the presence of a charge density wave, see Fig. 1(b).
@@ III.4 Properties of the anyon superconductor
At $2/3$ filling, both options we discussed are topologically ‘didd-id’ superconductors. Owing to the fact that the neutral spin modes on the boundary are counter-propagating and remain unmodified close to the critical point, the superconducting state retains a spin response $\frac{2}{4\pi}A_{s}\wedge\mathrm{d}A_{s}$ which is opposite to the electronic Hall effect. Due to the nature of the FQH parent states at general fillings $\nu=2/(2l+1)$, the sign of the spin-quantum Hall effect (and therefore also the thermal Hall response) alternates as $\sigma_{xy}^{s}=2\cdot(-1)^{l}$. Interestingly, while the spin-quantum Hall response is inherited from the parent states, the chiral central charge changes, i.e. from $c_{-}=0$ in the $2/3$ FQH state to $c_{-}=-2$ in the superconductor.
@@
As we are describing an insulator-to-insulator transition, this behavior is expected. Remarkably, since our doping procedure preserves the spin-gap, we expect that it also sets the size of Cooper pairs for weak doping. This implies Cooper pairs are inherently tightly bound close to the transition, while the topology is consistent with a chiral $d+id$ superconductor in the BCS limit. The tight binding of the Cooper pairs implies that the critical temperature $T_{c}$ is set by the stiffness $\rho_{s}$ of the superconductor, rather than the pairing-scale. Starting from each parent state, the stiffness is expected to grow with the doping level and suppressed by the anyon mass $m_{*}$ such that $\rho_{s}\sim\delta,\;m_{*}^{-1}$, see Ref. [6] for a derivation.
@@
Furthermore, we note that charge-2e superconductivity with opposing thermal Hall response can be obtained by considering Kohn-Luttinger mechanisms in the complementary weakly interacting limit where the parent state is a chiral metal [48, 64].
@@ IV Microscopic lattice model
where $t$ ($t^{\prime}$) is the (next-)nearest neighbor hopping, $U$ are interlayer interactions and $V$ nearest-neighbor interactions both within and between the layers. The Peierls phases are chosen to match the flux through a triangle $\Phi=B_{z}a^{2}\sqrt{3}/4$, where $a$ is the lattice constant. At $\Phi=\frac{\Phi_{0}}{2}$ flux piercing a triangle, and $t^{\prime}=t/4$ the system exhibits nearly flat particle-hole symmetric valence and conduction bands, with Chern numbers $C=1$ and C=1C=-1, respectively; see Fig. 4(a). Our model explicitly breaks time-reversal symmetry, and the Chern numbers for the two flavors $\sigma$ are equal. At finite filling and intermediate interaction strength, the system can thus realize various bilayer quantum Hall states. The Hubbard-Hofstadter model discussed here should be thought of as a controlled system, with potential experimental implementation in moiré heterostructures [34]. It serves as proof of principle that the considerations of the previous sections can be realized in a microscopic lattice model, but we expect our mechanism to be applicable more generally.
@@
First, we establish that the Halperin $(\bar{1}\bar{1}2)$ state is stabilized in the bilayer Hubbard-Hofstadter model Eq. (51) on the triangular lattice. In contrast to the $(110)$ state discussed in Refs. [34, 9], the $(\bar{1}\bar{1}2)$ state has a non-trivial topological order and has fascinating properties in its own right. As discussed in Sec. II.1, it can be understood in a parton language as spinons filling Chern bands with C=1C=-1 and chargons in a bosonic $1/2$ Laughlin state. We uniquely identify the $(\bar{1}\bar{1}2)$ state from its entanglement spectrum and by performing adiabatic flux insertion into the cylinder.
@@
$\pm\partial_{t}\varphi_{s}\partial_{y}\varphi_{s}+v_{s}\partial_{y}\varphi_{s}\partial_{y}\varphi_{s}]$
for $(mmn)$ states with $m=n\pm 1$ and velocities $v_{c/s}>0$. Note that for $m=n+1$, the two edge modes have the same chirality, while for m=n1m=n-1, they have opposite chirality, resulting in a non-chiral entanglement spectrum, assuming $n>0$. We find the edge Hamiltonian and total momentum [14]
@@
Using the above, the edge spectrum accessible through the entanglement spectrum can be understood, see Appendix E for details. There are three dispersing modes: A pure charge mode with $L_{s}=0$, a pure spin mode with $L_{c}=0$, and a mixed mode with $L_{c}\neq 0\neq L_{s}$. The counting of the entanglement spectrum in Fig. 5 agrees well with the expected edge spectrum for the $(\bar{1}\bar{1}2)$ state for which m=1,n=2m=1,\,n=2, as graphically illustrated.
@@
Increasing the nearest-neighbor interactions triggers a phase transition and tunes the system into a different insulator, which we will argue is nothing but the semion crystal. While the charge response of this phase is trivial, spin-pumping reveals a quantized spin-Hall response, which is identical to the spin-Hall response of the $(\bar{1}\bar{1}2)$ state; see Fig. 6(b). For general semion crystals formed at filling $\nu=2/(2n\pm 1)$, the quantum spin Hall response is alternating $Q_{s}=\pm 1$. This can be understood in this limit as lattice translations are broken, and the unit cell is enlarged. Within the enlarged unit cell, the electron and spinon density is integer $(n_{e}=1)$, while the flux per unit cell of the crystal is $\pm\Phi_{0}/2$ modulo $2\pi$. Thus, the spinon density in the crystal is now large enough to absorb all of the external flux by forming an IQH state at level $\pm 2$, while the chargons are in a Mott state with respect to the enlarged unit cell. As a concrete example, consider the 2/3 FCI, which transitions into the semion crystal, shown in the top left of the phase diagram of Figure 4(b). While triangles of the lattice enclose $\pi/2$ flux, the larger triangles defined by the charge ordered state have three times the area and enclose a flux of $3\pi/2\equiv-\pi/2$, which reverses the sign of the spin Hall conductance compared to the chiral spin liquid at integer filling.
Figure 7: Evidence for second-order phase transition in the Hubbard-Hofstadter model. (a) Order parameter $\Delta_{\text{CDW}}$ as a function of the interaction strength $V$ for fixed $U=4t$, $\chi=2048$ and $L_{y}=4$. Inset: Density correlations $C(k_{y}=0,x)=\langle\rho_{-k_{y}}(x)\rho_{k_{y}}(0)\rangle$ for $V=1.1t$ and $V=2.0t$. Due to the finite cylinder circumference, translations are already weakly broken in the FQH phase. (b) The correlation length of the density mode, i.e. of operators with zero charge and spin $(Q,S_{z})=(0,0)$, peaks at a certain interaction strength $V$ which we identify as the critical point (indicated by a dashed line in all panels). By contrast, the spin correlation length $(Q,S_{z})=(0,\hbar)$ remains small and does not grow when approaching the transition. (c) The ratio of $\xi_{2e}$ $(Q,S_{z})=(2,0)$] to $\xi_{1e}$ $(Q,S_{z})=(1,\hbar/2)$] correlation length is much larger than the trivially expected value of $0.5$ and peaks close to the critical point, growing as a function of the bond dimension $\chi$.